OpenStudy (anonymous):

PLEASE HELP> DUE TOMORROW Use the fundamental principles of counting to find the number of outcomes for the situation of flipping a fair coin six times. can you find the probability that you will get heads all six times? I started a chart but i hhave no idea what to do and my teacher doesnt explain

OpenStudy (goformit100):

@sami-21

OpenStudy (anonymous):

you have two choices for each toss, H and T

OpenStudy (anonymous):

if you flip 6 times, the counting principle tells you there are \[2\times2\times2\times2\times2\times2=2^6=64\] possible outcomes

OpenStudy (anonymous):

i know the possible outcome is 64 but idk how to get the other number :/

OpenStudy (anonymous):

that is, 64 possible combinations of heads and tails there is only one way to get all heads

OpenStudy (anonymous):

that would look like H, H, H, H , H, H

OpenStudy (anonymous):

so your answer is \[\frac{1}{64}\]

OpenStudy (anonymous):

but what about the other answers? idk how to get them

OpenStudy (anonymous):

it would be harder if the question was "what is the probability you get 2 heads and 4 tails, but this one was easier because there is only one way to do it

OpenStudy (anonymous):

what others? that is the only question i see posted

OpenStudy (anonymous):

lol im so confused hold on ill attatch a picture of my chart i had to draw

OpenStudy (anonymous):

this!

OpenStudy (anonymous):

ah now that is a different question

OpenStudy (anonymous):

whoops! can you tell my teacher hasnt taught me anything? its the first day of school too

OpenStudy (anonymous):

you need all the numerators, each one of which is \(\dbinom{n}{k}\) here \(n=6\) and \(k=0,1,2,3,4,5,6\)

OpenStudy (anonymous):

idk why that was so spaced out

OpenStudy (anonymous):

i am often spaced out we can do this an easy way

OpenStudy (anonymous):

so the n would fill out to be 6,12, and 18?

OpenStudy (anonymous):

real easy look at the 6th row of pascal's triangle and copy off the number for the numerator

OpenStudy (anonymous):

no you compute like this \[\dbinom{6}{1}=\dbinom{6}{5}=6\] \[\dbinom{6}{2}=\dbinom{6}{4}=\frac{6\times 5}{2}=3\times 5=15\] \[\dbinom{6}{3}=\frac{6\times 5\times 4}{\ 3\times 2}=5\times 4=20\] those are your numerators

OpenStudy (anonymous):

how did you decide to do the 6 ovver 1 6 over 5 6 over 2 and 6 over 4???

OpenStudy (anonymous):

it is called "n choose k" and it is the number of ways to choose \(k\) items out of \(n\) look at the 6th row of pascals triangle here http://ptri1.tripod.com/ you will see the numbers \[1,6,15,20,15,6,1\] and those are your numerators

OpenStudy (anonymous):

if you have not seen this, i have no idea what you are supposed to do, unless they want you to write down all 64 possible combinations of heads and tails and count

OpenStudy (anonymous):

haha yeah but thank you sooo much for helping me with the answer !! (: